OpenStudy (anonymous):

Four cards are randomly chosen (without replacement) from a deck of 52 cards. Let K be the event that two of the four cards are King and the other two are Aces. Let E be the event that one of the four cards is a King of Hearts. The conditional probability P(K | E) is equal to

OpenStudy (anonymous):

is the answer 0.01531? thanks!

OpenStudy (anonymous):

else, do let me know the steps leading to the solution! thanks!!

OpenStudy (kropot72):

The events K and E are independent, the reason being that the probability of K occurring is unaffected by whether or not E has occurred. In this case\[P(K|E)=P(K)\] P(K) can be found using the hypergeometric distribution.

OpenStudy (anonymous):

thanks! So, will this be correct then: P(K)=4!⋅4^2⋅3^2/(52⋅51⋅50⋅49⋅4) ?

OpenStudy (kropot72):

First I calculated the probability of getting 2 kings in a sample of 4 cards without replacement:\[P(2\ kings)=\frac{\left(\begin{matrix}4 \\ 2\end{matrix}\right)\left(\begin{matrix}48 \\ 2\end{matrix}\right)}{\left(\begin{matrix}52 \\ 4\end{matrix}\right)}=0.0249995\] The probability of getting 2 aces in a sample of 4 cards without replacement will be the same as P(2 kings). The events 2 kings and 2 aces are independent. therefore the probability of getting 2 kings and 2 aces is0.0249995^2 = 0.000625

OpenStudy (agent0smith):

@kropot72 didn't you forget to take into account that one of the kings has to be a heart?

OpenStudy (agent0smith):

I think they're also dependent events aren't they... we're finding the probability that we get 2 kings and 2 aces, given that one of the cards is the king of hearts.

OpenStudy (kropot72):

@agent0smith The question states that four cards are randomly chosen (without replacement). If you start looking at the selection and notice one of the cards is a King of Hearts does that affect the probability of two Kings and two Aces? I think not. That was the view taken in my attempt at a solution. However on reconsideration I accept your point in which case we get \[P(K|E)=\frac{P(K \cap E)}{P(E)}\]

OpenStudy (agent0smith):

But it's the probability that we get 2 K and 2 A, *given* one card is already a K of hearts. That should definitely affect the probability.

OpenStudy (agent0smith):

ie we can't have a K of spades and a K of clubs, or K diamonds and K clubs etc. Having one heart limits our possible outcomes, and thus makes K depend on E.

OpenStudy (kropot72):

I am not sure whether you agree or disagree with the conditional probability of K given E being expressed as\[P(K|E)=\frac{(K \cap E)}{P(E)}\] Note that the numerator on the right hand side is K intersection E which is the redefined sample space resulting from the consideration of the condition that a King of Hearts is included.

OpenStudy (agent0smith):

No I wasn't disagreeing with that, I think that's right. I was just describing why the events are dependent.

OpenStudy (kropot72):

Sure. I thought I covered that in saying previously that I accept your point :)

OpenStudy (agent0smith):

hehe, I just wasn't sure what you meant above, if you agreed or not :P